The center of gravity of a 45gram meter stick is at the 49 cm. mark. The stick is part of a lever which has the fulcrum at the 73 cm. mark. Where must an 86gram object be hung in order to have equilibrium? Show solutions.




1
Expert's answer
2022-04-26T14:56:06-0400

"M=Fl,"

"0.045\cdot(0.73-0.49)=0.086l,"

"l=12.5~cm,"

"x=73+12.5=85.5~cm."